# Heights and periods of algebraic cycles in families

### v1

@@ 4.3. A finiteness result for ${\mathscr{N}}^{0}$
Let $Y$ be a weakly optimal subvariety of $S$. Take a complex analytic irreducible component $\tilde{Y}$ of $u^{-1}(Y)$ such that $\tilde{Y}\cap\mathfrak{F}\not=\emptyset$. Then $\tilde{Y}$ is Zariski optimal in $u^{-1}(S)$ by Corollary 4.3. Then by Proposition 4.5, we have $\tilde{Y}^{\mathrm{ws}}=H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}^{0}$.
Hence by the finiteness result Proposition 4.6, there exist finitely many ${\mathbb{Q}}$-groups $N_{1},\ldots,N_{k}<\mathbf{G}$ satisfying the following property: For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}=u(\tilde{Y}^{\mathrm{ws}})$ equals $u(N_{j}({\mathbb{R}})^{+}x)$ for some $j\in\{1,\ldots,k\}$ and some $x\in{\mathscr{D}}$ with $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})$. Then ${\mathscr{D}}^{\prime}_{j}:=\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})^{+}x$ is a Mumford–Tate domain and is independent of the choice of such $x$. We are done. ∎
5. Application to non-degeneracy in case of irreducible VHSs
In this section, we give two applications of Theorem 3.1: when the VHS is irreducible and when it has a simple algebraic monodromy group. Both cases apply to the Gross-Shoen and the Ceresa normal functions.
@@ 5. Application to non-degeneracy in case of irreducible VHSs
We wish to check whether $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$, which by Corollary D.7 becomes $r(\nu)=\dim S$ for the Betti rank $r(\nu)$ defined in (1.4). Now Theorem 3.1 gives a checkable criterion. Recall the period map $\varphi=\varphi_{\nu}\colon S\rightarrow\Gamma\backslash{\mathscr{D}}$, the ${\mathbb{Q}}$-group $\mathbf{G}$ and its unipotent radical $V$ which is a vector group from (2.1). We emphasize that $V$ is in general not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$ and its geometric meaning is given by Remark 2.2: $\frac{1}{2}\dim V$ is the relative dimension of ${\mathscr{J}}_{\nu}$, the smallest intermediate Jacobian of sub-VHS of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ translated by a torsion multisection which contains $\nu(S)$.
@@
Set $\mathbf{G}_{0}=\mathbf{G}/V$. Then $V$ is a $\mathbf{G}_{0}$-submodule of ${\mathbb{V}}_{{\mathbb{Q}},s}$ for one (and hence all) $s\in S({\mathbb{C}})$.
By (i), ${\mathbb{V}}_{{\mathbb{Q}},s}$ is irreducible as a $\mathbf{G}_{0}$-module. By (ii), $V\not=\{0\}$. Hence $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
Let $N$ be a normal subgroup of $\mathbf{G}$. Then $V\cap N$ is a $\mathbf{G}_{0}$-submodule of $V$, and hence gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$. Hence by (i), either $V\cap N=\{0\}$ or $V\cap N=V$.
Assume $V\cap N=\{0\}$. Then $N\lhd\mathbf{G}_{0}=\mathbf{G}/V$ is reductive and it acts trivially on $V=V/(V\cap N)$. Hence, $N({\mathbb{R}})^{+}x_{0}$ is a point for any $x_{0}\in{\mathscr{D}}_{0}$, and therefore $N$ is contained in the center of $\mathbf{G}_{0}$. Using again the fact that $N$ acts trivially on $V$, we see that $N$ is contained in the center of $\mathbf{G}$. So $\dim\varphi_{/N}(S)=\dim\varphi(S)$ in this case.
On the other hand, it is clearly true that $\min_{N,~{}V\cap N=V}\left\{\dim\varphi_{/N}(S)\right\}+\frac{1}{2}\dim V$ is attained at $N=\mathbf{G}$ and hence equals $\frac{1}{2}\dim V$. Now we are done. ∎
@@
Set $\mathbf{G}_{0}=\mathbf{G}/V$. By Deligne, $H\lhd\mathbf{G}^{\mathrm{der}}_{0}$.
Let $N$ be a normal subgroup of $\mathbf{G}$. We may and do assume $N\lhd\mathbf{G}^{\mathrm{der}}$. The reductive part $\mathbf{G}_{N}:=N/(V\cap N)$ is a normal subgroup of $\mathbf{G}$. Now that $\mathbf{G}_{N}\cap H$ is a normal subgroup of $H$, we have either $\mathbf{G}_{N}\cap H=\{1\}$ or $H<\mathbf{G}_{N}$ by (i).
Assume $H<\mathbf{G}_{N}$. Since $\mathbf{G}_{0}$ is reductive, we can decompose $V=(V\cap N)\bigoplus(V\cap N)^{\perp}$ as $\mathbf{G}_{0}$-modules. Now $(V\cap N)^{\perp}$ gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$. But $H$ acts trivially on $(V\cap N)^{\perp}$ because $\mathbf{G}_{N}$ acts trivially on $V/(V\cap N)$. So $(V\cap N)^{\perp}=\{0\}$ by (ii). Hence $V\cap N=V$, and hence $V<N$. Thus $\varphi_{/N}(S)$ is a point. So $\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)=\frac{1}{2}\dim V$.
If $\mathbf{G}_{N}\cap H=\{1\}$, then $\dim\varphi_{/N}(S)=\dim\varphi(S)$. Now we are done. ∎
Now, we apply these two theorems to the Gross-Shoen and the Ceresa normal functions. Let $\nu_{{\mathrm{GS}}}$ (resp. $\nu_{{\mathrm{Ce}}}$) be the Gross–Schoen (resp. Ceresa) normal function ${\mathscr{M}}_{g}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in (C.5), with ${\mathbb{V}}_{{\mathbb{Z}}}$ from Proposition C.10. Let $\beta_{{\mathrm{GS}}}$ and $\beta_{{\mathrm{Ce}}}$ be the associated Betti forms.
@@
We shall apply Theorem 5.1 and let us check the assumptions. Assumption (i) holds by Proposition C.10. (iii). Assumption (ii) holds by Proposition C.10. (ii). Moreover the VHS ${\mathbb{V}}_{{\mathbb{Z}}}$ on $S$ has maximal moduli, so $\dim\varphi_{{\mathrm{GS}}}({\mathscr{M}}_{g})=\dim\varphi_{{\mathrm{Ce}}}({\mathscr{M}}_{g})=\dim{\mathscr{M}}_{g}=3g-3$ where $\varphi_{{\mathrm{GS}}}$ and $\varphi_{{\mathrm{Ce}}}$ are the period maps. Finally, $\dim{\mathbb{V}}_{{\mathbb{Q}},s}=\frac{2g(2g-1)(2g-2)}{6}-2g$, which is $>2(3g-3)$ when $g\geq 3$. Hence, we can conclude by Theorem 5.1. ∎
@@
Set $U:=S\setminus S_{{\mathscr{F}}}(1)$, which is Zariski open by Theorem 3.2.
By Corollary 5.3 and the “In particular” part of Theorem 3.2, we have that $\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}=0$. Thus $S_{{\mathscr{F}}}(1)\not=S$. Now, part (i) follows from Proposition C.7.
For (ii), notice that by definition $S_{{\mathscr{F}}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion. Thus, the set $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is discrete, and hence at most countable. ∎
6. Northcott property for Gross–Schoen and Ceresa cycles
@@ 6.1. Construction of the adelic line bundle
The following theorem is a reformulation of the second-named author’s [Zha10, Thm. 2.3.5].
@@
In some neighborhood $U$ of $s$ in ${\mathscr{M}}_{g}$, $\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})=\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is an element of $Z^{2}({\mathscr{C}}_{U}^{3})$, i.e. it is a $2$-cocycle of ${\mathscr{C}}_{U}^{3}$.
@@
$|t_{1}|\cap|t_{2}|\cap|t_{3}|=\emptyset,\qquad|\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})|\cap|(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})|=\emptyset,$
@@
where $\hat{t}_{1,s}$, $\hat{t}_{2,s}$, $\hat{t}_{3,s}$ are suitable arithmetic divisors on some model of ${\mathscr{C}}_{s}^{2}$ extending $t_{1,s}$, $t_{2,s}$, $t_{3,s}$. The right hand side is precisely $\frac{1}{[{\mathbb{Q}}(s):{\mathbb{Q}}]}\widehat{\deg}\overline{{\mathscr{L}}}|_{s}=h_{\overline{{\mathscr{L}}}}(s)$. The left hand side is $\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}$ because, by (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$. Hence, we are done. ∎

### v2

@@ 4.3. A finiteness result for ${\mathscr{N}}^{0}$
Let $Y$ be a weakly optimal subvariety of $S$. Take a complex analytic irreducible component $\widetilde{Y}$ of $u^{-1}(Y)$ such that $\widetilde{Y}\cap\mathfrak{F}\not=\emptyset$. Then $\widetilde{Y}$ is Zariski optimal in $u^{-1}(S)$ by Corollary 4.3. Then by Proposition 4.5, we have $\widetilde{Y}^{\mathrm{ws}}=H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}^{0}$.
Hence by the finiteness result Proposition 4.6, there exist finitely many ${\mathbb{Q}}$-groups $N_{1},\ldots,N_{k}<\mathbf{G}$ satisfying the following property: For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}=u(\widetilde{Y}^{\mathrm{ws}})$ equals $u(N_{j}({\mathbb{R}})^{+}x)$ for some $j\in\{1,\ldots,k\}$ and some $x\in{\mathscr{D}}$ with $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})$. Then ${\mathscr{D}}^{\prime}_{j}:=\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})^{+}x$ is a Mumford–Tate domain and is independent of the choice of such $x$. We are done. ∎
5. Application to non-degeneracy in some geometric cases
In this section, we give two applications of Theorem 3.1: when the VHS is irreducible or when it has a simple algebraic monodromy group. Both cases apply to the Gross-Shoen and the Ceresa normal functions.
@@ 5. Application to non-degeneracy in some geometric cases
We wish to check whether $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$, which by Corollary D.7 becomes $r(\nu)=\dim S$ for the Betti rank $r(\nu)$ defined in (1.5). Now Theorem 3.1 gives a checkable criterion. Recall the period map $\varphi=\varphi_{\nu}\colon S\rightarrow\Gamma\backslash{\mathscr{D}}$, the ${\mathbb{Q}}$-group $\mathbf{G}$ and its unipotent radical $V$ which is a vector group from (2.3). We emphasize that $V$ is, in general, not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$ and its geometric meaning is given by Remark 2.2: $\frac{1}{2}\dim V$ is the relative dimension of ${\mathscr{J}}_{\nu}$, the smallest intermediate Jacobian of sub-VHS ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ of ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow S$ translated by a torsion multisection which contains $\nu(S)$.
@@
We can say more in the situation of Theorem 5.1. In fact in this case ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}={\mathbb{V}}_{{\mathbb{Z}}}$. By the comment below (2.8), the equality $r(\nu)=\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}$ can be attained only if ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is an abelian scheme. Hence $\dim{\mathbb{V}}_{{\mathbb{Q}},s}>2\dim\varphi(S)$ in the situation of Theorem 5.1 if ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is not an abelian scheme.
@@
Set $\mathbf{G}_{0}=\mathbf{G}/V$. Then $V$ is a $\mathbf{G}_{0}$-submodule of ${\mathbb{V}}_{{\mathbb{Q}},s}$ for one (and hence all) $s\in S({\mathbb{C}})$, and $\mathbf{G}_{0}$ is by definition a subgroup of ${\mathrm{GL}}({\mathbb{V}}_{{\mathbb{Q}},s})$.
By (i), ${\mathbb{V}}_{{\mathbb{Q}},s}$ is irreducible as a $\mathbf{G}_{0}$-module. By (ii), $V\not=\{0\}$. Hence $V={\mathbb{V}}_{{\mathbb{Q}},s}$. Thus we have $\mathbf{G}_{0}\subset{\mathrm{GL}}(V)$. Let $N$ be a normal subgroup of $\mathbf{G}$. Then $V\cap N$ is a $\mathbf{G}_{0}$-submodule of $V$, and hence gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$. Hence by (i), either $V\cap N=\{0\}$ or $V\cap N=V$.
In the first case, we have the image $N_{0}$ of $N$ acting trivially on $V$. Thus $N_{0}=0$. This shows that $N$ is trivial. So $r(\nu)=\dim\varphi(S)$.
In the second case, $V\cap N=V$. It is clearly true that
$\min_{N,~{}V\cap N=V}\left\{\dim\varphi_{/N}(S)\right\}+\frac{1}{2}\dim V$
is attained at $N=\mathbf{G}$ and hence equals $\frac{1}{2}\dim V$. ∎
@@
An important case where assumption (i) is satisfied is when the Mumford–Tate group of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is quasi-simple; this follows from Deligne [Del87, §$\S$7.5]. And similar to the comment below Theorem 5.1, we must have $\dim V>2\dim\varphi(S)$ in the situation of Theorem 5.2 if ${\mathscr{J}}({\mathbb{V}}^{\prime}_{{\mathbb{Z}}})\rightarrow S$ is not an abelian scheme.
@@
Set $\mathbf{G}_{0}=\mathbf{G}/V$. By Deligne [Del87, §$\S$7.5], $H\lhd\mathbf{G}^{\mathrm{der}}_{0}$.
Let $N$ be a normal subgroup of $\mathbf{G}$. The reductive part $N_{0}:=N/(V\cap N)$ is a normal subgroup of $\mathbf{G}_{0}$. Now that $N_{0}\cap H$ is a normal subgroup of $H$, by (i), we have that either $N_{0}\cap H$ is finite or $H<N_{0}$.
Assume $H<N_{0}$. Since $\mathbf{G}_{0}$ is reductive, we can decompose $V=(V\cap N)\bigoplus(V\cap N)^{\perp}$ as $\mathbf{G}_{0}$-modules. Now $(V\cap N)^{\perp}$ gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$. But $H$ acts trivially on $(V\cap N)^{\perp}$ because $N_{0}$ acts trivially on $V/(V\cap N)$. So $(V\cap N)^{\perp}=\{0\}$ by (ii). Hence $V\cap N=V$, and hence $V<N$. So in this case we have
$\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)\geq\frac{1}{2}\dim V.$
The equality obtains when $N=\mathbf{G}$.
Assume that $N_{0}\cap H$ is finite. To handle this case, we again use the universal period map of the VHS $\widetilde{\varphi}_{0}\colon\widetilde{S}\rightarrow{\mathscr{D}}_{0}$ associated with the VHS ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow S$ from (2.3), where ${\mathscr{D}}_{0}$ is a $\mathbf{G}_{0}({\mathbb{R}})^{+}$-orbit and the natural projection $p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}$ is induced by the quotient $\mathbf{G}\rightarrow\mathbf{G}_{0}$. Moreover, we have $\widetilde{\varphi}_{0}=p\circ\widetilde{\varphi}$.
By logarithmic Ax [GK24, Thm. 7.2], we have $\widetilde{\varphi}_{0}(\widetilde{S})\subseteq H({\mathbb{R}})^{+}x_{0}$ for some $x_{0}\in{\mathscr{D}}_{0}$. So $\widetilde{\varphi}(\widetilde{S})\subseteq p^{-1}(\widetilde{\varphi}_{0}(\widetilde{S}))\subseteq p^{-1}(H({\mathbb{R}})^{+}x_{0})$. Notice that $p^{-1}(H({\mathbb{R}})^{+}x_{0})=(V\rtimes H)({\mathbb{R}})^{+}x$ for some $x\in{\mathscr{D}}$. So $\widetilde{\varphi}(\widetilde{S})\subseteq(V\rtimes H)({\mathbb{R}})^{+}x$, and hence $\varphi(S)\subseteq u((V\rtimes H)({\mathbb{R}})^{+}x)$ for $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$.
[…]
$p_{N}\colon\mathbf{G}\xrightarrow{p_{V\cap N}}\mathbf{G}/(V\cap N)\xrightarrow{p_{N_{0}}}\mathbf{G}/N,$
and this induces
$[p_{N}]\colon\Gamma\backslash{\mathscr{D}}\xrightarrow{[p_{V\cap N}]}\Gamma_{/V\cap N}\backslash({\mathscr{D}}/(V\cap N))\xrightarrow{[p_{N_{0}}]}\Gamma_{/N}\backslash({\mathscr{D}}/N).$
It is clearly true that $\dim\varphi(S)=\dim([p_{V\cap N}]\circ\varphi)(S)$. On the other hand, $\dim([p_{V\cap N}]\circ\varphi)(S)=\dim([p_{N_{0}}]\circ[p_{V\cap N}]\circ\varphi)(S)$ by the conclusion of the previous paragraph. So $\dim\varphi_{/N}(S)=\dim\varphi(S)$ in this case.
Now, we are done by combining the two cases above. ∎
Now, we apply these two theorems to the Gross–Schoen and the Ceresa normal functions. Let $\nu_{{\mathrm{GS}}}$ (resp. $\nu_{{\mathrm{Ce}}}$) be the Gross–Schoen (resp. Ceresa) normal function ${\mathscr{M}}_{g}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in (C.6), with ${\mathbb{V}}_{{\mathbb{Z}}}$ from Proposition C.10. Let $\beta_{{\mathrm{GS}}}$ and $\beta_{{\mathrm{Ce}}}$ be the associated Betti forms.
@@
We shall apply Theorem 5.1 and let us check the assumptions. Assumption (i) holds by part (iii) of Proposition C.10. Assumption (ii) holds by part (ii) of Proposition C.10. Moreover the VHS ${\mathbb{V}}_{{\mathbb{Z}}}$ on $S$ has maximal moduli, so $\dim\varphi_{{\mathrm{GS}}}({\mathscr{M}}_{g})=\dim\varphi_{{\mathrm{Ce}}}({\mathscr{M}}_{g})=\dim{\mathscr{M}}_{g}=3g-3$ where $\varphi_{{\mathrm{GS}}}$ and $\varphi_{{\mathrm{Ce}}}$ are the period maps. Finally, $\dim{\mathbb{V}}_{{\mathbb{Q}},s}=\frac{2g(2g-1)(2g-2)}{6}-2g$, which is $>2(3g-3)$ when $g\geq 3$. Hence, we can conclude by Theorem 5.1. ∎
@@
Set $U:=S\setminus S^{\mathrm{Betti}}(1)$, which is Zariski open by Theorem 3.2.
By Corollary 5.3 and the “In particular” part of Theorem 3.2, we have that $\min\{t\geq 0:S^{\mathrm{Betti}}(t)=S\}=0$. Thus $S^{\mathrm{Betti}}(1)\not=S$. Now, part (i) follows from Proposition C.7.
For (ii), notice that by definition $S^{\mathrm{Betti}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion. Thus, the set $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is discrete, and hence at most countable. ∎
6. Beilinson–Bloch height of Gross–Schoen and Ceresa cycles
@@ 6.1. Construction of the adelic line bundle
The line bundle $\overline{{\mathscr{L}}}$ has been constructed [Yua21, §$\S$3.3.4] in slightly different step.
The following theorem is a reformulation of the second-named author’s [Zha10, Thm. 2.3.5]. Recall the notation ${\mathrm{GS}}({\mathscr{C}}_{s}):={\mathrm{GS}}(\xi_{s})$ with $\xi_{s}\in\mathrm{Pic}^{1}({\mathscr{C}}_{s})$ such that $(2g-2)\xi_{s}$ being the canonical divisor class on ${\mathscr{C}}_{s}$.
@@
As notation, for any subset $U\subseteq{\mathscr{M}}_{g}$, set $\xi_{U}\in\mathrm{Pic}^{1}({\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}U/U)$ to be such that $(2g-2)\xi_{U}$ is the canonical divisor class of ${\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}U/U$. In some neighborhood $U$ of $s$ in ${\mathscr{M}}_{g}$, $\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})=\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$ is an element of $Z^{2}({\mathscr{C}}_{U}^{3})$, i.e. it is a $2$-cocycle of ${\mathscr{C}}_{U}^{3}$.
@@
$|t_{1}|\cap|t_{2}|\cap|t_{3}|=\emptyset,\qquad|\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})|\cap|(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})|=\emptyset,$
@@
where $\hat{t}_{1,s}$, $\hat{t}_{2,s}$, $\hat{t}_{3,s}$ are suitable arithmetic divisors on some model of ${\mathscr{C}}_{s}^{2}$ extending $t_{1,s}$, $t_{2,s}$, $t_{3,s}$. The right hand side is precisely $\frac{1}{[{\mathbb{Q}}(s):{\mathbb{Q}}]}\widehat{\deg}\overline{{\mathscr{L}}}|_{s}=h_{\overline{{\mathscr{L}}}}(s)$. The left hand side is $\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}$ because, by (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\xi_{U}}({\mathscr{C}}_{U})$. Hence, we are done. ∎